OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. –4 + yi = x + 3i Steps.........

hartnn (hartnn):

when we equate two complex numbers, their real and imaginary parts become equal. so if -4+yi = x+3i u get -4 = x and y = 3 so (x,y) = (-4,3) ok ?

hartnn (hartnn):

@sara1234 did u understand that ^^ ?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

is that all?

hartnn (hartnn):

yup, u got your x and y values :)

OpenStudy (anonymous):

are you sure the real numbers x and y

hartnn (hartnn):

-4 and 3 are real numbers only as they don't contain 'i' .

OpenStudy (anonymous):

thanks for your help really apperciate it :)

hartnn (hartnn):

welcome :)